OpenStudy (anonymous):

Please help. I will fan and give medal to whoever succeeds in helping me. For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (anonymous):

-y = 8-4x So when you switch the negative what happens?

OpenStudy (anonymous):

I'm not sure what you mean by "switch the negative".

OpenStudy (anonymous):

do the opposite sign so 8= -8

OpenStudy (anonymous):

When you divide out the negative what happens to (8-4x)

OpenStudy (anonymous):

so what is the opposite of -4 ?

OpenStudy (anonymous):

4.

OpenStudy (anonymous):

Yep @warningiflirt is right. -1 means the opposite. So take the opposite of 8-4x

OpenStudy (anonymous):

so you have −8 + 4x

OpenStudy (anonymous):

see how I did that?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

-y = (8 -4x) y = -(8-4x) If you know how to distribute it makes it y = -8 + 4x

OpenStudy (anonymous):

Cool